If the target bodyweight of a flock of sheep with BCS of 3.5 is 80 kg;

(i) Calculate the current bodyweight of the sheep if their present BCS is 2.5

(ii) If the sheep are expected to gain 180 grams per day on average, how long (days) will it take the animals to attain the target bodyweight of 80 kg

(iii) The ME requirement of sheep is 4.40 Mcal/day and DM intake is 2.1 kg/day; calculate the daily amount of supplement (ME = 3.0 Mcal/kg) and oat hay (ME = 1.99 Mcal/kg) that will be fed to the sheep
Answer #1

The BCS (Body condition score) of the sheep is usually represented in a five score scale ranging from 1 to 5.

About 12% of the current body weight is considered for a sheep to move one score up or down the given score.

It is given that a flock of sheep with BCS 3.5 has bodyweight 80kg.

Hence,

12% of 80kg = 0.12 * 80 = 9.60 kgs

9.60 kg is the weight representing on unit score increase or decrease in the sheep.

(i) If the flcok of sheep has BCS 2.5 which is 1 less than the target body weight score 3.5,

The current body weight can be given as:

80 - 9.6 = 70.4 kgs.  

(ii) If the sheep are expected to gain 180 grams per day on an average.

180 grams = 0.18 kgs

Number of days it would require for the sheep flock to achieve the target body weight 80 kg would be:

9.6 / 0.18 = 53.3 ~ 54 days.

(iii) ME requirement of sheep is 4.40 Mcal/day and DM intake is 2.1kg/day

Amount of supplement (ME = 3.0 Mcal/kg) and oat hay (ME = 1.99 Mcal/kg) that will be fed to the sheep would be:

Amount of supplement = 1 kg

Amount of oat hay = 1.1 kgs

Therefore, Energy gained = (1 * 3Mcal/kg) + (1.1 * 1.99Mcal/kg) = 5.189 Mcal/day which is greater than Minimum energy requirement